OpenStudy (anonymous):

Let X be normal with mean 80 and variance 9. Find P(X < 81) and P(78 < X < 82) Let X be normal with mean 80 and variance 9. Find P(X < 81) and P(78 < X < 82) @Mathematics

OpenStudy (anonymous):

P(x<81) = P (z<1/3) =.6293 from the z tables

OpenStudy (anonymous):

z= 81-80/3 =1/3

OpenStudy (anonymous):

P(78<x<82) = P(x<82) - P(x<78)

OpenStudy (anonymous):

P(x<82) = P(z<2/3) = 0.7486

OpenStudy (anonymous):

P(x<78)=P(Z<-2/3) = 1 - .7486 =.2514

OpenStudy (anonymous):

so P(78<x<82) = .7486-.2514 = .4972

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

I understand the last part, but for the first part, according to the book, it gives an answer of 0.6306. I think it may be wrong, but I am inclined to think it may be wrong

OpenStudy (anonymous):

hmm may be wrong on the book

OpenStudy (anonymous):

or they have estimated between the z values I used 1/3 = 0.33 they may have been more acurate using0.333333333333 etc

OpenStudy (anonymous):

or used the graphing calculator

OpenStudy (anonymous):

perhaps. it also says the value for the other problem is 0.4950. so they have gotten different values as well for that. i don't know

OpenStudy (anonymous):

using the Ti83 calculator and nomalcdf(-1E99, 81, 80, 3) = 0.6305

OpenStudy (zarkon):

It's just rounding error

OpenStudy (anonymous):

nomalcdf(78, 82, 80, 3) =0.4950

OpenStudy (anonymous):

so they must be using technology to get the answer and not the tables

OpenStudy (anonymous):

you may be right. thanks though, to the both of you, for your answer. at least i know my answer was not off.